One year​ ago, your company purchased a machine used in manufacturing for $ 95 comma 000. You have learned that a new machine is available that offers many​ advantages; you can purchase it for $ 160 comma 000 today. It will be depreciated on a​ straight-line basis over ten​ years, after which it has no salvage value. You expect that the new machine will contribute EBITDA​ (earnings before​ interest, taxes,​ depreciation, and​ amortization) of $ 40 comma 000 per year for the next ten years. The current machine is expected to produce EBITDA of $ 23 comma 000 per year. The current machine is being depreciated on a​ straight-line basis over a useful life of 11​ years, after which it will have no salvage​ value, so depreciation expense for the current machine is $ 8 comma 636 per year. All other expenses of the two machines are identical. The market value today of the current machine is $ 50 comma 000. Your​ company's tax rate is 35 %​, and the opportunity cost of capital for this type of equipment is 11 %. Is it profitable to replace the​ year-old machine? The NPV of the replacement is ​$ nothing. ​(Round to the nearest​ dollar.)
Answer #1

Answer:

No,

It is not profitable to replace the​ one year-old machine, since NPV is negative.

NPV of the replacement is = - $17019

Workings:

Old machine is purchased one year ago at = $95,000

Annual depreciation = $8636

Book value now = $95,000 - $8636 = $86,364

Market value of old machine = $50,000

Loss on sale = 86364 - 50000 = $36364

Sale value net of tax = 50000 + 36364 * 35% = $62,727.40

1 to 10 Year New machine cost Sale proceeds from old machine net of tax New Machine - EBITDA Old Machine - EBITDA Incremental